# The Narrow Depth and Breadth of Corporate Responsible AI Research

### v1

@@ 7.2 Direction of AI is on a Suboptimal Path
The direction of AI research has far-reaching consequences for power, politics, and social status in society (Acemoglu,, 2021). In particular, the path-dependent nature of technological progress (Arthur,, 2009) implies that decisions made during the design and implementation phases of AI will have profound societal implications once the technology is widely adopted (Bijker,, 1997; Winner,, 1980). The negative impact of AI systems can often be difficult to reverse (Ehsan et al.,, 2022), underscoring the importance of carefully considering potential consequences and guiding AI development toward socially beneficial outcomes from the very beginning.
@@
If market forces fail to steer AI innovation towards socially beneficial outcomes, proactive public policies and democratic processes are needed to mitigate potential harms and ensure that AI aligns with the greater good (Rosenberg,, 1969; Acemoglu,, 2021). Just as subsidies have been introduced to encourage innovation in renewable energy (Acemoglu et al.,, 2012), similar measures could be taken to address the challenges associated with AI and other digital technologies. As AI progresses, governments can provide guidance on which validated AI systems are appropriate to deploy in high-stakes public contexts, considering factors like efficacy, safety, fairness, and transparency. By establishing a robust policy framework now, we can guide the development of human-complementary AI in a direction that enhances human potential and benefits society as a whole (Brynjolfsson,, 2022). Carefully crafted policies are essential to ensure that AI progress aligns with societal values.
@@ 7.3 Limitations
Our study is not without limitations. First, our focus on peer-reviewed responsible AI publications may not capture the full extent of AI firms’ research engagement. While our research questions led us to examine only peer-reviewed responsible AI publications, AI firms may engage in research that they do not publish. Additionally, it is also possible that firms have papers that are not peer-reviewed. However, this lack of public engagement is a concern because prior research suggests that research engagement with academics also increases the capacity for firms to adopt responsible AI research (Baruffaldi and Poege,, 2023; Cohen and Levinthal,, 1990). Additionally, peer-reviewed research and engagement with academic outlets are important in building trust helping to work through important and ambiguous design choices (Ghoshal and Dasgupta,, 2023; Kleinberg et al.,, 2023).
Second, since the diffusion of knowledge from basic science to product creation requires time (Mansfield,, 1991), our patent citation analysis may underestimate the actual extent of the commercialization of responsible AI research. Moreover, it is important to recognize that not all elements of responsible AI research are amenable to patenting, nor are they always reflected in patent citations (Cohen et al.,, 2000). However, we observe that some AI patents cite recent academic papers on conventional AI and non-AI, suggesting that industry can, to some extent, build upon academic work in this area.
Third, the training data for our supervised machine learning model, sourced from leading responsible AI conferences, might not fully represent all industry research activities. While these conferences address a wide range of issues, they may still miss some industry research topics. To address this, we implemented measures such as data ablation, using different subsets of training datasets to classify industry papers (see Fig. S4 in the appendix), and conducted manual validation. Additionally, we validated our findings using a different classification method–an expert-consulted keyword search (see S7)–which supported our key results. Lastly, we used a comprehensive sample of papers identified through keyword searches (presented in Fig. 2c), yielding similar outcomes.
Notwithstanding these limitations, our results suggest that industry has a limited breadth and depth in responsible AI research and rarely cites research in its commercial innovation.
@@ 8 Conclusion
AI is set to transform our society, but like all technologies, it also carries notable risks (Acemoglu,, 2021; Lazar,, 2022; Acemoglu and Lensman,, 2023; Brynjolfsson,, 2022). Our large-scale study shows that the organizations that are developing and deploying this technology have limited public engagement in responsible AI research. More specifically, we document that even within AI research firms, who are setting the research agenda and driving innovation (Ahmed et al.,, 2023; Frank et al.,, 2019; Koch et al.,, 2021), only 11.2% engage in meaningful responsible AI research. Additionally, we find that firms demonstrate a limited scope in their approach to responsible AI and show a lack of enthusiasm toward critical issues in this area. Our results are robust across different datasets and measurements. This study emphasizes that industry’s participation in responsible AI research is markedly limited in breadth and depth. This limited participation is despite industry’s substantial resources and significant external pressure (The White House, 2023a, ) to engage in this field. Additionally, our large-scale patent analysis shows a disconnect between research and commercialization, with industry largely failing to integrate responsible AI research into their commercial products and services. Our results underscore the urgent need for industry to prioritize and deepen its involvement in responsible AI research; ensuring that the development and deployment of AI align with societal values and mitigate potential harms. Failure to do so may jeopardize public trust, hinder the realization of AI’s benefits, and exacerbate the risks posed by this powerful technology.
@@ Acknowledgements
Process for constructing Fig. 1b and 1c Refer to caption
@@
Figure S7: The first figure (Fig. S7a) shows the diagram of 1,771 AI firms that have AI research and responsible AI research activities using an expert-recommended keyword list. We present them in three groups based on their degree of engagement: (a) no engagement in responsible AI research, (b) limited engagement (published fewer than 5 papers), and (c) engagement (published 5 or more papers). Fig. S7b showcases a stark difference in the leading 100 AI firms’ (n = 506,017 papers) and universities’ (n = 5,265,419 papers) participation in AI research compared to their responsible AI research. Here, individual organizations are symbolized by discrete dots. A trend line reflects the participation trend within each group, and the shaded bands correspond to 95% confidence intervals. Additionally, the dashed line indicates a reference line where the proportion of responsible AI papers to total AI papers is 2.5%.
@@
S1.4 Keywords for Classifying AI Research Papers
@@
S1.5 Industry’s limited engagement in responsible AI research at conventional AI conferences
We examined industry engagement in responsible AI at conventional AI conferences. In recent years, conventional AI conferences have also started to focus on responsible AI research (Maslej et al.,, 2024). It is plausible that industry is doing more work in responsible AI but not presenting them at the leading responsible AI conferences.
@@
Industry has a narrower focus on responsible AI research than academia: We used publication data from the three leading responsible AI conferences and used the abstracts of those papers to identify key research themes in responsible AI research. Fig. S9 shows the research topics that industry and academia explore in their responsible AI research. We used k-means clustering analysis on the abstracts of the responsible AI papers (see keywords list in Table S8 in the appendix). After clustering, we labeled each paper based on the cluster and then calculated the percentage of papers associated with that cluster for each year. Here, we present the analysis of 10 clusters on papers from 2018 to 2022 (the results are similar for other clusters, see Fig. S12 for 8 clusters & S13 for 6 clusters). Our analysis shows that relative to academia, industry engages in fewer topics (only 4 topics in 2018 compared to academia’s 8 topics). The tendency for industry to have a more limited research scope than academia, as shown across other years, aligns with prior research (Klinger et al.,, 2020). It suggests a consistent pattern of industry’s narrower research focus over time. Moreover, our findings remain consistent even when we employ alternative machine learning methodologies like LDA topic modeling (see Fig. S10) and bigram frequency count (see Fig. S14), or theme-specific analysis (see Fig. 3b).
@@
Furthermore, we found that industry typically began exploring research topics only after academia introduced them. For instance, we observed that industry’s four topics in 2018 increased to seven by the end of 2022 (see Fig. S9). Conversely, academia started with a broader spectrum of topics (8 topics), which eventually increased to 10. Interestingly, we observed that industry research has broadened more over time, underscoring the importance of industry’s consistent engagement in academic venues (Baruffaldi and Poege,, 2023). This analysis provides suggestive evidence that research engagement can benefit industry by broadening its understanding of responsible AI. Besides, the expansion of industry’s research portfolio was, in part, likely influenced by the relevance and utility of those additional topics to them. Overall, AI firms have a narrow scope of responsible AI research and appear to trail behind academia in exploring a range of topics within this field.
@@ 10 Topic- 0: fairness, individuals, ai, ethics, show, different, social, ethical, machine, research
Figure S20: This figure shows industry and academic papers’ cumulative citations in USPTO patents. By matching against a comprehensive patent-paper citation data (n = 32,698,465; 1947-2022), we show in Fig. S20a and S20c that, only 88 and 396 responsible AI papers from industry and academia, respectively, have been cited in generic patents, while 15,236 academia and 7,532 industry conventional AI and non-AI papers have been cited between 2010-22. Using a separate dataset of AI patents (n = 141,770; 1985-2018), Fig. S20b and S20d further documents, 3 industry and 17 academic responsible AI papers have been cited in any AI patents between 2010-22, in stark contrast to the incorporation of conventional AI and non-AI research.
S5 Additional Details
Table S10: Recent news on firings/layoffs of responsible AI teams in Industry
Date Title Firms Quotes
Nov 14, 2022 Why Meta and Twitter’s AI and ML layoffs matter | The AI Beat | VentureBeat,
Elon Musk Has Fired Twitter’s “Ethical AI” Team | WIRED
Twitter, Meta “Ten days ago, as part of mass Twitter layoffs, the company’s entire ethical artificial intelligence (AI) team–which worked to make Twitter’s algorithms more transparent and fair–was let go. ”
“Meta layoffs an entire 50-person research team focused on machine learning (ML) infrastructure, called Probability.”
Mar 14, 2023 Microsoft lays off AI ethics and society team-The Verge Microsoft “The ethics and society team was at its largest in 2020, when it had roughly 30 employees including engineers, designers, and philosophers. In October, the team was cut to roughly seven people as part of a reorganization.”
Mar 29, 2023 Big tech companies cut AI ethics staff, raising safety concerns | Financial Times “The speed with which they are being abolished leaves Big Tech’s algorithms at the mercy of advertising imperatives, undermining the wellbeing of kids, vulnerable people and our democracy.”- Josh Simons, former Facebook AI ethics researcher and author of Algorithms for the People.
Mar 30, 2023 Microsoft and Amazon have cut their AI ethics teams as Elon Musk and Steve Wozniak call for a moratorium on tech companies’ “out-of-control race” Twitter, Microsoft “Twitter’s ‘ethical AI team’ was cut days before the first round of Musk’s layoffs affecting thousands of employees at the company in November, less than a week after he became CEO.”
Microsoft, “But it no longer has a central team of ethicists dedicated to researching the potential harms of AI systems. Members of the original AI ethics team were either reassigned to product teams or were laid off.”
April 12, 2023 Amazon’s Twitch Safety, AI Ethics Job Cuts Raise Concerns Among Ex-Workers - Bloomberg Amazon “Amazon’s Twitch Safety, AI Ethics Job Cuts Raise Concerns Among Ex-Workers”
May 26, 2023 Tech layoffs ravage the teams that fight online misinformation and hate speech Meta, Amazon, Alphabet (Google), Twitter “Meta, Amazon, Alphabet, and Twitter have all drastically reduced the size of their teams focused on internet trust and safety as well as ethics as the companies focus on cost cuts.”
“As part of Meta’s mass layoffs, the company ended a fact-checking project that had taken half a year to build, according to people familiar with the matter.”
May 30, 2023 Big Tech’s Cost-Cutting Slashes AI Ethics, Trust And Safety; Will Netflix’s Anti-Password Sharing Work? | AdExchanger Meta, Twitter, Amazon, Microsoft “Meta also laid off 200 content moderators, 16 members of Instagram’s well-being group, and more than 100 employees working on platform integrity.”
“Twitter slashed its ethical AI team from 17 to just one member and laid off 15% of its trust and safety department.”
“Amazon also downsized its ethical AI team and cut 50 positions dedicated to identifying abusive and illegal behavior on its streaming platform Twitch.”
“Microsoft cut all 30 members of its ethics and society team.”

### v2

@@ 7.2 Implications for the Direction of AI Development
The direction of AI research has implications for power, politics, and social dynamics in society (Acemoglu, 2021). The path-dependent nature of technological progress (Arthur, 2009) implies that decisions made during the design and implementation phases of AI will have enduring societal implications once the technology is widely adopted (Bijker, 1997; Winner, 1980). The effects of AI systems can be difficult to reverse (Ehsan, Singh, Metcalf, and Riedl, 2022), highlighting the value of considering potential consequences and guiding AI development toward beneficial outcomes during early stages.
@@
To the extent that market forces do not fully direct AI innovation toward socially beneficial outcomes, public policies and democratic processes may play a role in addressing potential challenges and supporting alignment between AI and broader societal goals (Rosenberg, 1969; Acemoglu, 2021). Just as subsidies have been introduced to encourage innovation in renewable energy (Acemoglu, Aghion, Bursztyn, and Hemous, 2012), similar measures could be considered to address challenges associated with AI and other digital technologies. As AI progresses, governments can provide guidance on which validated AI systems are appropriate to deploy in high-stakes public contexts, considering factors like efficacy, safety, fairness, and transparency. By establishing policy frameworks, societies can guide the development of human-complementary AI in directions that enhance human potential and benefit society (Brynjolfsson, 2022). Thoughtfully designed policies can support alignment between AI progress and societal values.
@@ 7.3 Limitations
Our study is not without limitations. First, our focus on peer-reviewed responsible AI publications may not capture the full extent of AI firms’ research engagement. While our research questions led us to examine only peer-reviewed responsible AI publications, AI firms may engage in research that they do not publish. Additionally, firms may have papers that are not peer-reviewed. That said, prior research suggests that research engagement with academics increases firms’ capacity to adopt responsible AI research (Baruffaldi and Poege, 2023; Cohen and Levinthal, 1990). Additionally, peer-reviewed research and engagement with academic outlets are important for building trust and working through important and ambiguous design choices (Ghoshal and Dasgupta, 2023; Kleinberg, Mullainathan, and Raghavan, 2023).
Second, since the diffusion of knowledge from basic science to product creation requires time (Mansfield, 1991), our patent citation analysis may underestimate the actual extent to which responsible AI research informs commercial innovation. Moreover, not all elements of responsible AI research are amenable to patenting, nor are they always reflected in patent citations (Cohen, Nelson, and Walsh, 2000). However, we observe that some AI patents cite recent academic papers on conventional AI and non-AI, suggesting that industry can, to some extent, build upon academic work in this area.
Third, the training data for our supervised machine learning model, sourced from leading responsible AI conferences, might not fully represent all industry research activities. While these conferences address a wide range of issues, they may still miss some industry research topics. To address this, we implemented measures such as data ablation, using different subsets of training datasets to classify industry papers (see Fig. S4 in the appendix), and conducted manual validation. Additionally, we validated our findings using two alternative classification methods, an expert-consulted keyword search (see S7) and LLM-based classification (see S8), both of which produced consistent results.
Notwithstanding these limitations, our findings document patterns of industry engagement in responsible AI research and provide an empirical baseline for future studies examining the relationship between research activity and commercial innovation in this field.
@@ 8 Conclusion
AI is poised to transform society, and like all technologies, its development trajectory carries important implications (Acemoglu, 2021; Lazar, 2022; Acemoglu and Lensman, 2023; Brynjolfsson, 2022). Our large-scale study documents patterns of industry engagement in responsible AI research. We find that among AI research firms, which are shaping the research agenda and driving innovation (Ahmed, Wahed, and Thompson, 2023; Frank, Wang, Cebrian, and Rahwan, 2019; Koch, Denton, Hanna, and Foster, 2021), 34.5% publish responsible AI papers, and that the ratio of responsible AI to conventional AI research is lower in industry than in academia. We also observe that industry’s responsible AI research concentrates in a narrower range of topics compared to academic research, with greater emphasis on technical dimensions such as explainability and algorithmic fairness. These patterns are consistent across different datasets and measurements. Our large-scale patent analysis further documents that industry patents cite responsible AI research at lower rates than conventional AI research, though firms with active responsible AI research programs are also the most frequent citers. These findings provide an empirical baseline for understanding how responsible AI research is distributed across sectors and integrated into commercial innovation. As AI continues to advance, ongoing measurement of these patterns may inform discussions among researchers, practitioners, and policymakers regarding the development and deployment of AI systems.
@@ Acknowledgements
Process for constructing Fig. 1a and 1b Refer to caption
@@
Figure S7: Fig. S7 shows the engagement of 1,771 AI firms that have AI research and responsible AI research activities using an expert-recommended keyword list. It showcases a heterogeneity in the leading 100 AI firms’ (n = 506,017 papers) and universities’ (n = 5,265,419 papers) participation in AI research compared to their responsible AI research. Here, individual organizations are symbolized by discrete dots. A trend line reflects the participation trend within each group, and the shaded bands correspond to 95% confidence intervals. Additionally, the dashed line indicates a reference line where the proportion of responsible AI papers to total AI papers is 2.5%.
S1.4 Validation with LLM-based Classification
Industry’s engagement in responsible AI research (a comparison with academia)
Figure S8: This figure shows the engagement of AI firms in responsible AI research using LLM-based classification with in-context learning. Fig. S8 presents the distribution of the leading 100 AI firms’ (n = 506,017 papers) and universities’ (n = 5,265,419 papers) participation in AI research compared to their responsible AI research. Here, individual organizations are symbolized by discrete dots. A trend line reflects the participation trend within each group, and the shaded bands correspond to 95% confidence intervals. Additionally, the dashed line indicates a reference line where the proportion of responsible AI papers to total AI papers is 2.5%.
@@
S1.5 Keywords for Classifying AI Research Papers
@@
S1.6 Industry’s limited engagement in responsible AI research at conventional AI conferences
We examined industry engagement in responsible AI at conventional AI conferences. In recent years, conventional AI conferences have also started to focus on responsible AI research (Maslej, Fattorini, Perrault, Parli, Reuel, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Niebles, Shoham, Wald, and Clark, 2024). It is plausible that industry is doing more work in responsible AI but not presenting them at the leading responsible AI conferences.
@@
Industry has a narrower focus on responsible AI research than academia: We used publication data from the three leading responsible AI conferences and used the abstracts of those papers to identify key research themes in responsible AI research. Fig. S10 shows the research topics that industry and academia explore in their responsible AI research. We used k-means clustering analysis on the abstracts of the responsible AI papers (see keywords list in Table S8 in the appendix). After clustering, we labeled each paper based on the cluster and then calculated the percentage of papers associated with that cluster for each year. Here, we present the analysis of 10 clusters on papers from 2018 to 2022 (the results are similar for other clusters, see Fig. S13 for 8 clusters & S14 for 6 clusters). Our analysis shows that relative to academia, industry engages in fewer topics (only 4 topics in 2018 compared to academia’s 8 topics). The tendency for industry to have a more limited research scope than academia, as shown across other years, aligns with prior research (Klinger, Mateos-Garcia, and Stathoulopoulos, 2020). It suggests a consistent pattern of industry’s narrower research focus over time. Moreover, our findings remain consistent even when we employ alternative machine learning methodologies like LDA topic modeling (see Fig. S11) and bigram frequency count (see Fig. S15), or theme-specific analysis (see Fig. 3b).
@@
Furthermore, we found that industry typically began exploring research topics only after academia introduced them. For instance, we observed that industry’s four topics in 2018 increased to seven by the end of 2022 (see Fig. S10). Conversely, academia started with a broader spectrum of topics (8 topics), which eventually increased to 10. Interestingly, we observed that industry research has broadened more over time, underscoring the importance of industry’s consistent engagement in academic venues (Baruffaldi and Poege, 2023). This analysis provides suggestive evidence that research engagement can benefit industry by broadening its understanding of responsible AI. Besides, the expansion of industry’s research portfolio was, in part, likely influenced by the relevance and utility of those additional topics to them. Overall, AI firms have a narrow scope of responsible AI research and appear to trail behind academia in exploring a range of topics within this field.
@@ 10 Topic- 0: fairness, individuals, ai, ethics, show, different, social, ethical, machine, research
Figure S21: This figure shows industry and academic papers’ cumulative citations in USPTO patents. By matching against a comprehensive patent-paper citation data (n = 32,698,465; 1947-2022), we show in Fig. S21a and S21c that, only 88 and 396 responsible AI papers from industry and academia, respectively, have been cited in generic patents, while 15,236 academia and 7,532 industry conventional AI and non-AI papers have been cited between 2010-22. Using a separate dataset of AI patents (n = 141,770; 1985-2018), Fig. S21b and S21d further documents, 3 industry and 17 academic responsible AI papers have been cited in any AI patents between 2010-22, in stark contrast to the incorporation of conventional AI and non-AI research.